In the laboratory you dissolve 13.9 g of potassium hydroxide in a volumetric flask and add water to a total volume of 375 . mL.


What is the molarity of the solution?


1
Expert's answer
2021-02-22T05:52:35-0500

no. Of moles of potassium hydroxide = 13.9/56.1056 = 0.247 moles

Volume = 375ml = 0.375L


Molarity = no. Of moles/ volume

Molarity = 0.247/0.375 = 0.658 M